OpenStudy (anonymous):

A radiator contains 160 ounces of a 20% antifreeze solution. How much must be drained and replaced with pure antifreeze to increase the concentration of the solution 50%.

OpenStudy (anonymous):

well nice question!! first we calculate how much ounce of antifreeze substance is there in radiator... so in 160 ounce of 20% solution out of 160 ounce we have \[160 \times 20/100=32 \] so 32 ounce antifreeze substance is dissolved in solution so weight of remaining solution is 160-32=128 ounce now we want to increase concentration of antifreeze solution to 50% so what we we do we drain 32 ounce of solution so now we have 128-32=96 ounce of solution out of this 96 ounce we replace 32 ounce with antifreeze substance (replacing means subtract 32 ounce from weight of remaining solution and add 32 ounce to antifreeze substance) so now we have 64 ounce of antifreeze substance + 64 ounce of solution=128 ounce total now take % \[100 \times 64/128 =50 \] that's it i hope this helps..! :)

OpenStudy (anonymous):

@believegrl43